OpenStudy (anonymous):

decide whether the pair of lines is parallel, perpendicular, or neither 5x+3y=6 5x+3y+9

OpenStudy (anonymous):

parallel

OpenStudy (anonymous):

Parallel, if I assume correctly that you mean \[l: z=5x+3y+6\]and \[k: z=5x+3y+9\] The two lines will each travel in the same direction, they will only start at a different position on the z-axis. With these particular two equations, line k would be 3 points higher on the z-axis than line l.

OpenStudy (anonymous):

i think you're assuming incorrectly

OpenStudy (anonymous):

he stated no z's therefore he is referring to a system in R2 not R3

OpenStudy (anonymous):

she's*

OpenStudy (anonymous):

plus her previous questions imply that her grade level is between 8 and 10, not university or calculus 2 or 3. just sayin

OpenStudy (anonymous):

i thiink she meant 5x+3y=9

OpenStudy (anonymous):

you're trying to correct the first equation to mean 5x+3y+6

OpenStudy (anonymous):

instead

OpenStudy (anonymous):

thanks for the attack matt, Im 4 months from finishing my associates degree with a 3.79 GPA. Unfortunately I am terrible at algebra, everbody has weaknesses.

OpenStudy (anonymous):

and yes it was = not + ust a typo

OpenStudy (anonymous):

just*

OpenStudy (anonymous):

Then I stand corrected, and am not sure whether it is indeed parallel, perpendicular or none.